OpenStudy (anonymous):

The length of one edge of this cube is 4 cm. What is the surface area of the cube?

OpenStudy (michele_laino):

we have to apply this formula: \[A = 6 \times {L^2}\] where L is the edge of the cube, and A is the requested area

OpenStudy (anonymous):

I don't get it

OpenStudy (michele_laino):

hint: \[\large A = 6 \times {L^2} = 6 \times {4^2} = ...?\]

OpenStudy (anonymous):

24?

OpenStudy (anonymous):

24cm squared

OpenStudy (michele_laino):

no, it is: \[\large A = 6 \times {L^2} = 6 \times {4^2} = 6 \times 16 = 96\;c{m^2}\]